# Fixed-Time Input-to-State Stability for Singularly Perturbed Systems via Composite Lyapunov Functions

### v3

@@ 4 Fixed-Time Feedback Optimization with Time-varying Costs
In this section, we leverage the results of Theorems 1-3 to study a practical problem of interest in the context of singular perturbations: feedback optimization under slowly-varying cost functions [26]. In contrast to the existing asymptotic results [27, 28, 32, 33], we introduce an optimization-based controller able to achieve FxT stability via non-Lipschitz feedback. In particular, we consider plants of the form
@@
Assumption 4 is the “fixed-time” version of standard open-loop stability assumptions considered in the setting of feedback optimization [7, 27]. In particular, by taking y=zh(x^)y=z-h($\hat${x}) to quantify the deviation of $z$ from its steady-state approximation, the conditions of Assumption 4 simply ask that such deviation converges to zero by a fixed-time, for each fixed $\hat{x}\in\mathbb{R}^{n}$. For a general class of linear and nonlinear plants, this property can be achieved using different types of non-smooth controllers that combine super-linear and sub-linear feedback [34], homogeneity tools [35], or the implicit Lyapunov technique [36], to name just a few examples.
Our primary goal is to design a control law on the input $\hat{x}$ that stabilizes (25) in a fixed time the solution of the following time-varying optimization problem
@@ 4.1 Fixed-Time Gradient-Based Feedback
and hence, for each $\theta$, when the plant dynamics (25) are negligible, the dynamics (32) converge to the solution of (30) in fixed time [15, 16]. Since in our setting the dynamics (25) cannot be neglected, we can obtain a real-time feedback controller by replacing h(x^)h($\hat${x}) in (32) with the measured value of $z$ to obtain the following closed-loop system:
@@
We study system (34) under the following mild Lipschitz assumption, which is standard in the literature [27]:
@@
$~~~+c|x|^{2}+\frac{M^{2}}{c}|\varepsilon_{0}\Pi(\theta)|^{2}$
$\leq-\frac{\kappa^{2-\xi_{1}}}{L}V^{1-\frac{1}{2}\xi_{1}}{(x)}-\frac{\kappa^{2-\xi_{2}}}{L}V^{1-\frac{1}{2}\xi_{2}}{(x)}+\frac{M^{2}}{c}|\varepsilon_{0}\Pi(\theta)|^{2},$
where $0<c<\frac{1}{L}\min\left\{{\kappa^{2-\xi_{1}}},{\kappa^{2-\xi_{2}}}\right\}$, which shows that Assumption 2 is satisfied. Next, let $y=z-h(x+\varphi(\theta))$, which leads to the following dynamics:
@@
$~~~+\left\lvert\frac{P_{\theta}(x,y)}{|P_{\theta}(x,y)|^{\xi_{2}}}-\frac{P_{\theta}(x,0)}{|P_{\theta}(x,0)|^{\xi_{2}}}\right\rvert\bigg{)}$
$\leq 2^{\xi_{1}+1}\ell^{1-\xi_{1}}|x||y|^{1-\xi_{1}}$
$~~~+2K|x||y|\left(\ell^{-\xi_{2}}|y|^{-\xi_{2}}+L^{-\xi_{2}}|x|^{-\xi_{2}}\right)$
$\leq c\left(|x|^{2-\xi_{1}}+|x|^{2-\xi_{2}}\right)+\frac{1}{\mu_{1}(c)}\left(|y|^{2-\xi_{1}}+|y|^{2-\xi_{2}}\right)$
$\leq c\tilde{V}^{2}(x)+\frac{c_{1}^{\frac{1}{2}\xi_{1}-1}+c_{1}^{\frac{1}{2}\xi_{2}-1}}{\mu_{1}(c)}\tilde{W}^{2}(y),$
for all $c>0$, and $\mu_{1}\in\mathcal{K}_{\infty}$ is given by $\mu_{1}(c):=m^{-1}\left(c^{-\sigma_{1}}+c^{-\sigma_{2}}+c^{-\frac{1}{\sigma_{2}}}\right)^{-1}$, with $\sigma_{i}=\frac{1}{1-\xi_{i}}$ and $m:=\max\{(2^{\xi_{1}+1}\ell^{1-\xi_{1}})^{1+\sigma_{1}},(4K\ell^{-\xi_{2}})^{1+\sigma_{2}},(4K\ell^{-\xi_{2}})^{1+\frac{1}{\sigma_{2}}}\}$. Picking $c<\frac{1}{L}\min\left\{{\kappa^{2-\xi_{1}}},{\kappa^{2-\xi_{2}}}\right\}$ establishes (15a) and (15c). To verify (15b), we have
@@
Let $h^{*}:=\sup_{x}|\mathbf{J}_{h}(x)|$, which exists and is finite since $h$ is $\mathcal{C}^{1}$ and globally Lipschitz. We can proceed from (40) as follows:
@@
$\leq\tilde{L}\left(\tilde{V}^{2}+2\tilde{W}^{2}\right),$
where $\tilde{L}:=\eta h^{*}\max\{\ell^{1-\xi_{1}},L^{1-\xi_{1}},2^{-\xi_{2}}\ell^{1-\xi_{2}},2^{-\xi_{2}}L^{1-\xi_{2}}\}$. By applying Corollary 2 we obtain the result.

### v5

@@ 4 Fixed-Time Feedback Optimization with Time-varying Costs
In this section, we leverage the results of Theorems 1-3 to study a practical problem of interest in the context of singular perturbations: feedback optimization under slowly-varying cost functions [colombino2019online]. In contrast to the existing asymptotic results [hauswirth2020timescale, bianchin2022online, 9540998, 10189107], we introduce an optimization-based controller able to achieve closed-loop FxT stability via non-Lipschitz feedback. In particular, we consider plants of the form
@@
Assumption 4 is the “fixed-time” version of standard open-loop stability assumptions considered in the setting of feedback optimization [khalil, hauswirth2020timescale]. In particular, by taking y=zh(x^)y=z-h($\hat${x}) to quantify the deviation of $z$ from its steady-state approximation, the conditions of Assumption 4 simply ask that such deviation converges to zero by a fixed-time, for each fixed $\hat{x}\in\mathbb{R}^{n}$. For a general class of linear and nonlinear plants, this property can be achieved using different types of non-smooth controllers that combine super-linear and sub-linear feedback [andrieu2008homogeneous], homogeneity tools [10704051], or the implicit Lyapunov technique [POLYAKOV2015332], to name just a few examples.
Our primary goal is to design a control law on the input $\hat{x}$ that stabilizes (32), in a fixed time, at the solution of the following time-varying optimization problem
@@ 4.1 Fixed-Time Gradient-Based Feedback
and hence, for each $\theta$, when the plant dynamics (32) are negligible, the dynamics (39) converge to the solution of (37) in fixed time [fixed_time, nonsmoothesc]. Since in our setting the dynamics (32) cannot be neglected, we can obtain a real-time feedback controller by replacing h(x^)h($\hat${x}) in (39) with the measured value of $z$ to obtain the following closed-loop system:
@@
We study system (41) under the following mild Lipschitz assumption, which is standard in the literature [hauswirth2020timescale]:
@@
$~~~+\tilde{c}|x|^{2}+\frac{M^{2}}{\tilde{c}}|\varepsilon_{0}\Pi(\theta)|^{2}$
$\leq-\frac{\kappa^{2-\xi_{1}}}{L}V^{1-\frac{1}{2}\xi_{1}}{(x)}-\frac{\kappa^{2-\xi_{2}}}{L}V^{1-\frac{1}{2}\xi_{2}}{(x)}+\frac{M^{2}}{\tilde{c}}|\varepsilon_{0}\Pi(\theta)|^{2},$
where $0<\tilde{c}<\frac{1}{L}\min\left\{{\kappa^{2-\xi_{1}}},{\kappa^{2-\xi_{2}}}\right\}$, which shows that Assumption 2 is satisfied. Next, let $y=z-h(x+\varphi(\theta))$, which leads to the following dynamics:
@@
[…]
We denote $\Delta_{\theta}(x,y)=P_{\theta}(x,y)-P_{\theta}(x,0)$. Recall that from Assumption 7, we have $|\Delta_{\theta}(x,y)|\leq\ell|y|$. Hence, we can use Lemma 2 to obtain
$\left\lvert\frac{P_{\theta}(x,y)}{|P_{\theta}(x,y)|^{\xi_{1}}}-\frac{P_{\theta}(x,0)}{|P_{\theta}(x,0)|^{\xi_{1}}}\right\rvert$ $\leq 2^{\xi_{1}}|\Delta_{\theta}(x,y)|^{1-\xi_{1}}$
$\leq 2^{\xi_{1}}\ell^{1-\xi_{1}}|y|^{1-\xi_{1}}.$ (47)
We can again leverage Lemma 2 to obtain
$\left\lvert\frac{P_{\theta}(x,y)}{|P_{\theta}(x,y)|^{\xi_{2}}}-\frac{P_{\theta}(x,0)}{|P_{\theta}(x,0)|^{\xi_{2}}}\right\rvert$
$\leq K|\Delta_{\theta}(x,y)|\left(|P_{\theta}(x,0)|^{-\xi_{2}}+|\Delta_{\theta}(x,y)|^{-\xi_{2}}\right)$
$\leq K\ell|y|\left(L^{-\xi_{2}}|x|^{-\xi_{2}}+\ell^{-\xi_{2}}|y|^{-\xi_{2}}\right),$ (48)
where the bottom inequality uses the fact that $|P_{\theta}(x,0)|=|\nabla\Phi(x+\varphi_{\theta}(\theta))|\leq L|x|$. We can combine (47) and (48) to obtain
$|I_{1}|$ $\leq 2^{\xi_{1}+1}\ell^{1-\xi_{1}}|x||y|^{1-\xi_{1}}$
$~~~+{\color[rgb]{0,0,0}2K\ell}|x||y|\left(\ell^{-\xi_{2}}|y|^{-\xi_{2}}+L^{-\xi_{2}}|x|^{-\xi_{2}}\right)$ (49)
By Lemma 1, we have
$\delta_{1}|x||y|^{1-\xi_{1}}$ $\leq{c}|x|^{2-\xi_{1}}+\delta_{1}^{1+\sigma_{1}}{c}^{-\sigma_{1}}|y|^{2-\xi_{1}}$ (50a)
$\delta_{2}|x||y|^{1-\xi_{2}}$ $\leq{c}|x|^{2-\xi_{2}}+\delta_{2}^{1+\sigma_{2}}{c}^{-\sigma_{2}}|y|^{2-\xi_{2}}$ (50b)
$\delta_{3}|x|^{1-\xi_{2}}|y|$ $\leq{c}|x|^{2-\xi_{2}}+\delta_{3}^{1+\frac{1}{\sigma_{2}}}{c}^{-\frac{1}{\sigma_{2}}}|y|^{2-\xi_{2}}$ (50c)
for all c>0{c}>0, where $\sigma_{i}=\frac{1}{1-\xi_{i}}$, $\delta_{1}=2^{\xi_{1}+1}\ell^{1-\xi_{1}},\delta_{2}=2K\ell^{1-\xi_{2}}$, and $\delta_{3}=2K\ell L^{-\xi_{2}}$. Let $\tilde{V}(x)=|x|^{1-\frac{1}{2}\xi_{1}}+|x|^{1-\frac{1}{2}\xi_{2}}$ and $\tilde{W}(y)=W^{\frac{1}{2}b_{1}}(y)+W^{\frac{1}{2}b_{2}}(y)$. Since
$2b_{1}\leq 2-\xi_{1}\leq 2-\xi_{i}\leq 2-\xi_{2}\leq 2b_{2}$ (51)
holds for $i=1,2$, we also have that
$|x|^{2-\xi_{i}}$ $\leq|x|^{2-\xi_{1}}+|x|^{2-\xi_{2}}\leq\tilde{V}^{2}(x)$ (52a)
$|y|^{2-\xi_{i}}$ $\leq c_{1}^{\frac{1}{2}\xi_{i}-1}W^{1-\frac{1}{2}\xi_{1}}(y)\leq c_{1}^{\frac{1}{2}\xi_{i}-1}\tilde{W}^{2}(y)$ (52b)
for $i=1,2$.
We can then combine (49), (50), (52) to obtain
$|I_{1}|$ $\leq c(|x|^{2-\xi_{1}}+2|x|^{2-\xi_{2}})+\delta_{1}^{1+\sigma_{1}}{c}^{-\sigma_{1}}|y|^{2-\xi_{1}}$
$~~~+\delta_{2}^{1+\sigma_{2}}{c}^{-\sigma_{2}}|y|^{2-\xi_{2}}+\delta_{3}^{1+\frac{1}{\sigma_{2}}}{c}^{-\frac{1}{\sigma_{2}}}|y|^{2-\xi_{2}}$
Next, using
$m$ $:=\max\{\delta_{1}^{1+\sigma_{1}},\delta_{2}^{1+\sigma_{2}},\delta_{3}^{1+\frac{1}{\sigma_{2}}}\}$
$=\max\{(2^{\xi_{1}+1}\ell^{1-\xi_{1}})^{1+\sigma_{1}},{\color[rgb]{0,0,0}(2K\ell^{1-\xi_{2}})^{1+\sigma_{2}}},{\color[rgb]{0,0,0}(2K\ell L^{-\xi_{2}})^{1+\frac{1}{\sigma_{2}}}}\}$
we obtain:
$|I_{1}|$ $\leq c(|x|^{2-\xi_{1}}+2|x|^{2-\xi_{2}})$
$~~~~+m\Big({c}^{-\sigma_{1}}|y|^{2-\xi_{1}}+{c}^{-\sigma_{2}}|y|^{2-\xi_{2}}+{c}^{-\frac{1}{\sigma_{2}}}|y|^{2-\xi_{2}}\Big).$
Using ${\mu}_{1}(c)=\left(m\cdot\max\{c^{-\sigma_{1}},c^{-\sigma_{2}},c^{-\frac{1}{\sigma_{2}}}\}\right)^{-1}$ we obtain:
$|I_{1}|\leq c(|x|^{2-\xi_{1}}+{\color[rgb]{0,0,0}2}|x|^{2-\xi_{2}})+\frac{1}{{\mu}_{1}(c)}\left(|y|^{2-\xi_{1}}+{\color[rgb]{0,0,0}2}|y|^{2-\xi_{2}}\right).$
Therefore, using the definition of $\tilde{V}$, and by (52b), we have
$|I_{1}|$ $\leq{\color[rgb]{0,0,0}2}c\tilde{V}(x)+\frac{1}{{\mu}_{1}(c)}\left(c_{1}^{\frac{1}{2}\xi_{1}-1}\tilde{W}^{2}(y)+2c_{1}^{\frac{1}{2}\xi_{2}-1}\tilde{W}^{2}(y)\right)$
$={\color[rgb]{0,0,0}2}c\tilde{V}^{2}(x)+\frac{(c_{1}^{\frac{1}{2}\xi_{1}-1}+{\color[rgb]{0,0,0}2}c_{1}^{\frac{1}{2}\xi_{2}-1})}{\mu_{1}(c)}\tilde{W}(y)^{2}.$
for all $c>0$. Picking c{c} such that $0<c<\frac{1}{2L}\min\left\{{\kappa^{2-\xi_{1}}},{\kappa^{2-\xi_{2}}}\right\}$ establishes (15a) and (15c). To verify (15b), we have
@@
Let $h^{*}:=\sup_{x}|\mathbf{J}_{h}(x)|$, which exists and is finite since $h$ is $\mathcal{C}^{1}$ and globally Lipschitz. Moreover, we make note of the following fact:
$|P_{\theta}(x,y)|\leq|P_{\theta}(x,y)-P_{\theta}(x,0)|+|P_{\theta}(x,0)|\leq\ell|y|+L|x|.$
With this in mind, We can proceed from (53) as follows:
@@
$\leq\tilde{L}\left(\tilde{V}^{2}(x)+2\tilde{W}^{2}(y)\right),$
where $\tilde{L}:=\eta h^{*}\max\{\ell^{1-\xi_{1}}+L^{1-\xi_{1}},2^{-\xi_{2}}\left(\ell^{1-\xi_{2}}+L^{1-\xi_{2}}\right)\}$. By applying Corollary 2 we obtain the result.